# Exploring Multi-Document Information Consolidation for Scientific Sentiment Summarization

### v1

@@ Abstract
Modern natural language generation systems with LLMs exhibit the capability to generate a plausible summary of multiple documents; however, it is uncertain if models truly possess the ability of information consolidation to generate summaries, especially on those source documents with opinionated information. To make scientific sentiment summarization more grounded, we hypothesize that in peer review human meta-reviewers follow a three-layer framework of sentiment consolidation to write meta-reviews and it represents the logic of summarizing scientific sentiments in meta-review generation. The framework is validated via human annotation. Based on the framework, we propose evaluation metrics to assess the quality of generated meta-reviews, and we find that the hypothesis of the sentiment consolidation framework works out empirically when we incorporate it as prompts for LLMs to generate meta-reviews in extensive experiments. [^1] We will release the code and annotated data at ANONYMIZED upon acceptance.
@@ 1 Introduction
Notable strides have been made in abstractive text summarization (El-Kassas et al., 2021) with the advancement of large language models (LLMs) (Zhao et al., 2023) over recent years. With even a piece of simple instruction as a prompt such as “tl;dr” or “please write a summary”, these models can generate plausible summaries which are found more preferred over those written by humans (Pu et al., 2023). However, it is uncertain if these models truly possess the ability of information consolidation, especially when summarizing documents which are composed of opinionated information. The models may take shortcuts to generate texts instead of correctly understanding and aggregating information from the source documents (Gehrmann et al., 2023) and they may generate abstractive summaries with incorrect sentiment consolidation.
@@
Automated sentiment summarization holds significant importance (Kim et al., 2011) and there have been sentiment summarization datasets; however, most of them are in the product review domain. These datasets are not suitable for investigating information consolidation as (1) the summaries are synthetic, composed of a simple combination of extracted snippets (Amplayo et al., 2021), (2) the summary of product reviews is about extracting the majority sentiment. To address this, in this paper, we propose the task of scientific sentiment summarization, taking the meta-reviews in scientific peer review as summaries. [^2] The representative peer review platform which is publicly available is www.openreview.com. The investigation of meta-review generation (Li et al., 2023a) presents an exceptional opportunity for exploring the intricate process of multi-document information consolidation. This is because (1) meta-reviewers are supposed to understand not only all the reviews from different reviewers but also the multi-turn discussions between the reviewers and the author and write their comments to support the acceptance decision of the manuscript, (2) the argument logic has to be taken into account to arrive at sentiments in the meta-reviews where majority voting is not always working, (3) meta-reviews have to recognize and resolve conflicts and consensus among source documents (i.e., corresponding reviews and discussions).
In this paper, we hypothesize that human meta-reviewers are following a three-layer sentiment consolidation framework as shown in Figure 1 to write meta-reviews based on individual reviews and discussions in the peer review process. Human and automatic annotation is then conducted to get judgements from meta-reviews and corresponding source documents and these judgements are the foundation of our framework. With the proposed sentiment consolidation framework and annotated data, we propose two evaluation metrics which are focused more on sentiments in generated meta-reviews, and in our extensive experiments, we observe empirical validation of the framework hypothesis when integrated as prompts for LLMs to generate meta-reviews.
Contributions of the paper are summarized as follows:
@@
    With the initialization of scientific sentiment summarization, we hypothesize that human meta-reviewers follow a three-layer framework that describes the sentiment consolidation process in the meta-reviewing process;
@@
    We collect human annotations on meta-reviews and corresponding source documents based on the sentiment consolidation framework;
@@
    Experiments show that the framework works out empirically when we incorporate it as prompts for LLMs to generate meta-reviews.
@@
Content Expression
What the sentiment is talking about
Criteria Facet
The specific criteria facet that the judgement belongs to
Sentiment Polarity
The polarity and strength of the sentiment
Convincingness
How well the sentiment is justified in the document
Sentiment Expression
The value of the sentiment
@@
Min Max Average #Documents/Sample 5 30 12.4 #Words/Sample 1,541 11,901 4,260.9 #Words/Source document 10 1,562 360.53 #Words/Meta-review 16 648 150.87
@@ 2.1 Large-Scale Information Consolidation
Natural language generation systems are expected to not only have high-quality generations but also have the ability to comprehend the input information, especially for conditional text generation such as multi-document summarization which has to integrate and aggregate information from different source documents. Most work in the text summarization community is only trying to improve the generation quality of text summarization, such as relevance and faithfulness without considering the generation process. For example, Li et al. (2023b) use heterogeneous graphs to represent source documents and borrow the idea of graph compression to train the summarization model to get improvement of the generated summaries. However, it is uncertain if these models truly possess the ability to consolidate information from different source documents, and this is still under exploration. Moreover, it is also of limited exploration that the human information consolidation logic in summarization could be used to enhance the generation ability of LLMs and make generations more grounded.
@@ 2.2 Automated Sentiment Summarization
Sentiment summarization is to get a summary of sentiments in the input texts (Hossain et al., 2023). However, most datasets for sentiment summarization are in the product review domain (Amplayo et al., 2021), and scientific sentiment summarization is under exploration. Meta-review generation, which is a typical scenario of scientific sentiment summarization, is to automatically generate meta-reviews based on reviews and the multi-turn discussions between reviewers and the author of the corresponding manuscript (Li et al., 2023a). It is mostly modelled as an end-to-end task (Bhatia et al., 2020; Wu et al., 2022). Although Li et al. (2023a) take consideration of the in-nature conversational structure of source documents, their models have no knowledge of how human meta-reviewers write the meta-reviews and they do not consider the judgements which would be the most important content in the peer-reviewing process in their modelling or evaluation. Because this task features multiple source documents with complex relationships which should be considered in the corresponding meta-review, it is a suitable scenario to investigate the ability of models to consolidate information in the summarization process.
@@ 3 Sentiment Consolidation Over Multiple Opinionated Documents
In the following section, we introduce the task of scientific sentiment summarization and our three-layer sentiment consolidation framework in meta-review generation, conduct judgement identification and extraction, analyze the fusion process of scientific sentiments.
@@ 3.1 Hierarchical Sentiment Consolidation
We propose the task of scientific sentiment summarization, synthesizing reviews and discussions in the peer-reviewing process into the corresponding meta-review. While we use the PeerSum [^3] https://github.com/oaimli/PeerSum dataset where the input is reviews and discussions and the target output is the corresponding human-written meta-review, scientific sentiment summarization would focus more on sentiments in the summary or meta-review generation process.
In the domain of peer review, based on reviewer guidelines from the most popular academic presses such as ACM and IEEE [^4] The complete table of official guidelines that we consider is in the Appendix., we find that all comments in the whole process of peer-reviewing are mostly about judgements from different participants on the quality and merit of the manuscript. These opinions or judgements are based on six facets of review criteria (Novelty, Soundness, Clarity, Advancement, Compliance and Overall quality) which are the foundations of the peer-reviewing process. The meta-reviewers must form their final judgements based on those from the reviewers and authors. Similarly, based on meta-reviewer guidelines for ICLR [^5] https://iclr.cc/Conferences/2024/SACguide and NeurIPS [^6] https://nips.cc/Conferences/2020/PaperInformation/AC-SACGuidelines, we find that the meta-reviewer needs to understand and aggregate information from comments in the whole peer-reviewing process. They may first identify judgements from different review and discussion documents, then aggregate opinions in different criteria facets, and lastly based on this aggregation they write the meta-review.
To conceptualize the logic of scientific sentiment summarization from multiple documents, we propose a three-layer framework, taking meta-review generation as an example shown in Figure 1. The three layers include the input layer, the consolidation layer, and the generation layer. The input layer is to understand individual judgements in input documents of different types: official reviews and multi-turn discussions. The consolidation layer is to represent what the meta-reviewers are supposed to do to aggregate sentiments from the input documents. Specifically, they first identify and extract judgements from different documents, then reorganize the judgements, and lastly do judgement fusion to form the final opinions of each criteria facet. In the generation layer, the meta-reviewer writes the meta-review to express the judgements that they develop based on consolidation of the whole peer-reviewing process.
@@ 3.2 Judgement Identification and Extraction
Judgements lay the foundation of our proposed framework and the whole peer review process. A judgement here expresses sentiment on a criteria facet with sometimes its justification, and it contains several components: Content Expression, Sentiment Expression, Criteria Facet, Sentiment Level, and Convincingness Level (definitions are shown in Table 1, and examples are in Figure 5 in human annotation). To automate judgement identification and extraction, we first conduct human annotation, and then leverage in-context learning of LLMs to perform more (automatic) annotation.
@@
In human annotation, there are three types of documents including meta-reviews, official reviews, and discussions (the same definition to Li et al. (2023a)) to be annotated. We recruit two human annotators [^7] The two annotators are senior PhD students who are familiar with the peer-review process. to do this annotation. The whole human annotation instruction is in Appendix B. As the annotation is time-consuming, we finish 30 samples and it costs about one hour to annotate per sample and it costs about 60 hours and 2,100 US dollars in total. Two annotators get 1,812 and 1,744 judgements, respectively. The statistics of these 30 samples are present in Table 2. Agreement of the two annotators on these three types of documents is shown in Figure 2 in terms of Krippendorff’s $\alpha$ [^8] Calculation details and more results in terms of both Cohen’s $\kappa$ and Krippendorff’s $\alpha$ are in Table 10, Table 11 and Table 12 in Appendix C. It is clear that there is high agreement between the two human annotators. This shows that judgements on these criteria facets in these documents really exist and this makes it possible to analyze the sentiment consolidation among source documents.
@@
Figure 2: Inter-annotator agreement on meta-reviews, official reviews and discussions in terms of Krippendorff’s $\alpha$ for different judgement components including Content Expression (CE), Sentiment Expression (SE), Criteria Facet (CF), Sentiment Level (SL), and Convincingness Level (CL).
@@
Figure 3: The averaged GPT-4’s agreement with two human annotators on meta-reviews, official reviews and discussions in terms of Krippendorff’s $\alpha$ for different judgement components including Content Expression (CE), Sentiment Expression (SE), Criteria Facet (CF), Sentiment Level (SL), and Convincingness Level (CL).
To get more annotated judgements, we split the annotation task into two phases: extracting content and sentiment expressions, and predicting other components of judgements, and leverage GPT-4 to accomplish this with in-context learning as there is not enough data to train specific models. Specifically, we design prompts for GPT-4999The version of GPT-4 in the paper is gpt-4-0613 in default., which is currently one of the most powerful LLMs, with few-shot prompting to get the content and sentiment expressions which express the judgements in annotation documents (the prompt template is in Appendix D), and get the predictions of other judgement components (the prompt template is in  Appendix E). We present average agreement of GPT-4 with the two human annotators in Figure 3101010More results are in Table 13, Table 14 and Table 15 in Appendix C.. We can see that annotation results by GPT-4 have a high agreement with human annotation results, which means that GPT-4 can be qualified to do the annotation on meta-reviews and official reviews. It is also clear that GPT-4 cannot have high agreement with human annotators on annotating discussions. This may be because that judgements in discussion documents are usually to rebuttal and not easy to detect, which is consistent in human annotation. Moreover, it is clear that GPT-4 has lowest agreement with human annotators on convincingness levels in Figure 3 while human annotators have the lowest agreement on convincingness levels in Figure 2. This may be because that convincingness is subjective.
@@ 3.3 Sentiment Fusion for Consolidation
Table 3: Frequency of different criteria facets in meta-review judgements and meta-review documents.
With all the annotated judgements, we next dive more into the process of sentiment aggregation. Among all the criteria facets, we find that Soundness and Advancement are the two most important criteria facets when the meta-reviewers write their meta-reviews, while Compliance is rarely an issue in meta-reviews (shown in Table 3). This is consistent with our understanding of the peer-reviewing process.
More importantly, we find that human-written meta-reviews are not always following simple majority voting to develop sentiments. Specifically, in the document level, we find that in PeerSum there are 23.7% samples where human meta-reviewer’s acceptance decision is not consistent with the prediction based on majority voting by review ratings (if the number of ratings which are higher than 5 for a paper is larger than the number of ratings which are not higher than 5, the majority voting will get Accept while the human meta-reviewer’s decision may be Reject for the paper.); in the sentiment level, there is an example in Table 4 where the meta-reviewer is not following majority voting to form the sentiment on Novelty.
@@
To learn the aggregation function of sentiments from source documents to the meta-review, we try in-context learning of LLMs as there are not enough training data to train other models and LLMs like GPT-4 are powerful in zero-shot learning. We formulate this as a text classification task where the output is the sentiment level of the piece of Content Expression in a meta-review judgement and the input is judgements in all source documents which are belonging to the same criteria facet as the meta-review judgement or the full text of all source documents. Specifically, we randomly sample 100 human-annotated meta-review judgements for each criteria facet in each annotator’s annotation results, and the prompt for GPT-4 is in Appendix F. Average results between two human annotators in Table 5 suggest that using judgements is better than full texts to predict the meta-review sentiments and this zero-shot classification function has the potential to predict the sentiments in the meta-review. This suggests that it may be useful to incorporate some intermediate steps to extract text segments related to a facet before generating the final sentiment (i.e. these results argue that we should not treat it as a simple text-to-text problem where the input is reviews and discussions and the output is the meta-review).
Criteria Facets Judgements Full Texts Advancement 0.677 0.697 Soundness 0.684 0.667 Novelty 0.700 0.650 Overall 0.643 0.631 Clarity 0.712 0.645 Compliance 0.555 0.593
Table 5: Accuracy of GPT-4 to predict sentiment levels of human-annotated meta-review judgements based on annotated source judgements or original full texts.

### v2

# A Sentiment Consolidation Framework for Meta-Review Generation

@@ Abstract
Modern natural language generation systems with Large Language Models (LLMs) exhibit the capability to generate a plausible summary of multiple documents; however, it is uncertain if they truly possess the capability of information consolidation to generate summaries, especially on documents with opinionated information. We focus on meta-review generation, a form of sentiment summarisation for the scientific domain. To make scientific sentiment summarization more grounded, we hypothesize that human meta-reviewers follow a three-layer framework of sentiment consolidation to write meta-reviews. Based on the framework, we propose novel prompting methods for LLMs to generate meta-reviews and evaluation metrics to assess the quality of generated meta-reviews. Our framework is validated empirically as we find that prompting LLMs based on the framework — compared with prompting them with simple instructions — generates better meta-reviews. [^1] The code and annotated data are accessible at https://github.com/oaimli/MetaReviewingLogic.
@@ 1 Introduction
Notable strides have been made in abstractive text summarization (El-Kassas et al., 2021) with the advancement of Large Language Models (LLMs) (Zhao et al., 2023) over recent years. With even a simple instruction such as “tl;dr” or “please write a summary”, these models can generate plausible summaries which are found more preferred over those written by humans (Pu et al., 2023). However, it is uncertain if these models truly possess the ability of information consolidation, especially when summarizing documents that are composed of opinionated information. The models may take shortcuts to generate texts instead of correctly understanding and aggregating information from the source documents (Gehrmann et al., 2023) and they may generate abstractive summaries with incorrect overall sentiment.
@@
Automated sentiment summarization holds significant importance (Kim et al., 2011) and there have been sentiment summarization datasets; however, most of them are in the product review domain. These datasets are less interesting for investigating information consolidation as (1) the summaries are extractive, composed of a simple combination of extracted snippets (Amplayo et al., 2021), and (2) the summary of product reviews is about extracting the majority sentiment (which is a simple consolidation function). To address this, in this paper, we propose the task of scientific sentiment summarization, taking the meta-reviews in scientific peer review as summaries. [^2] The representative peer review platform which is publicly available is www.openreview.com. The investigation of meta-review generation (Li et al., 2023a) presents an exciting opportunity for exploring the intricate process of multi-document information consolidation that involves complex judgement. This is because (1) meta-reviewers are supposed to understand not only all the reviews from different reviewers but also the multi-turn discussions between the reviewers and the author and write their comments to support the acceptance decision of the manuscript, (2) the logic of arguments (from reviewers and authors) has to be taken into account to arrive at the final sentiment in the meta-reviews and it is not a matter of majority voting and (3) meta-reviews have to recognize and resolve conflicts and consensus among reviewers.
In this paper, we hypothesize that human meta-reviewers follow a three-layer sentiment consolidation framework as shown in Figure 1 to write meta-reviews based on reviews and multi-turn discussions in the peer review process. Human and automatic annotation is then conducted to extract sentiments and expressions on various review facets (e.g., novelty and soundness) from corresponding source documents (i.e., reviews and discussions) and these judgements play a critical role in generating the meta-reviews. We also propose two evaluation metrics which focus on assessing sentiments in generated meta-reviews, and experiments empirically validate our proposed three-layer framework when they are integrated as prompts for LLMs to generate meta-reviews.
Contributions of our paper:
@@
    We hypothesize that human meta-reviewers follow a three-layer sentiment consolidation framework when writing meta-reviews;
@@
    We collect human annotations on meta-reviews and corresponding source documents based on the consolidation framework;
@@
    Experiments validate the empirical effectiveness of the framework when we incorporate it as prompts for LLMs to generate meta-reviews.
@@
Content Expression What the sentiment is talking about
Sentiment Expression The value of the sentiment
Review Facet The specific review facet that the judgement belongs to
Sentiment Level The polarity and strength of the sentiment
Convincingness Level How well the sentiment is justified in the document
@@
Min Max Average #Documents/Sample 5 30 12.4 #Words/Sample 1,541 11,901 4,260.9 #Words/Source document 10 1,562 360.5 #Words/Meta-review 16 648 150.9
@@ 2.1 Large-Scale Information Consolidation
Natural language generation systems are expected to not only have high-quality generations but also have the ability to comprehend the input information, especially for conditional text generation such as multi-document summarization which has to integrate and aggregate information from different source documents (Gehrmann et al., 2023). Most work in the text summarization community only attempts to improve the generation quality of text summarization, such as relevance and faithfulness, without considering the intricate generation process (Phang et al., 2022; El-Kassas et al., 2021; Xiao et al., 2022). For example, Li et al. (2023b) use heterogeneous graphs to represent source documents and borrow the idea of graph compression to train the summarization model to get improvement of the generated summaries. However, it is uncertain if these models truly possess the ability to consolidate information from different source documents.
@@ 2.2 Automated Sentiment Summarization
Sentiment summarization aims to summarise the overall sentiment given a set of documents (Hossain et al., 2023). However, most datasets for sentiment summarization are in the product review domain (Amplayo et al., 2021), and scientific sentiment summarization is under-explored. Meta-review generation, which is a typical scenario of scientific sentiment summarization, is to automatically generate meta-reviews based on reviews and the multi-turn discussions between reviewers and the author of the corresponding manuscript (Li et al., 2023a). It is mostly modelled as an end-to-end task (Bhatia et al., 2020; Wu et al., 2022; Shen et al., 2022; Chan et al., 2020). Although Li et al. (2023a) considered the conversational structure of reviews and discussions, their models do not explain how human meta-reviewers write the meta-reviews. Different from investigating checklist-guided iterative introspection for meta-review generation with prompting (Zeng et al., 2024), our work is based on a three-layer sentiment consolidation framework and focuses on various review facets, and we explicitly investigate the sentiment fusion process which is arguably an important aspect of meta-review generation.
@@ 3 Sentiment Consolidation Over Multiple Opinionated Documents
In the following section, we introduce the task of scientific sentiment summarization and our three-layer sentiment consolidation framework in meta-review generation, conduct sentiment and expression extraction, and analyze the fusion process of scientific sentiments.
@@ 3.1 Hierarchical Sentiment Consolidation
The task is meta-review generation. We use the PeerSum [^3] https://github.com/oaimli/PeerSum dataset where the input is reviews and discussions and the target output is the corresponding human-written meta-review. We should clarify that even though the task is to generate meta-reviews, our focus here is to get the overall sentiment in the meta-reviews to be correct. Our method and evaluation reflect this focus.
Reading the reviewer guidelines from popular academic presses such as ACM and IEEE [^4] The complete table of official guidelines that we consider is in Appendix A., we find they are mostly about judgements on the quality and merit of the manuscript. These judgements are generally based on six review facets of criteria: Novelty, Soundness, Clarity, Advancement, Compliance and Overall quality. The meta-reviewers must form their final opinion based on these judgements from the reviewers and authors. Looking at the meta-reviewer guidelines for ICLR [^5] https://iclr.cc/Conferences/2024/SACguide and NeurIPS [^6] https://nips.cc/Conferences/2020/PaperInformation/AC-SACGuidelines, it recommends the meta-reviewer to understand and aggregate information from the whole peer-reviewing process. That is, a human meta-reviewer should first identify judgements from reviews and discussions, and then consolidate these opinions from different review facets to write their meta-review.
To conceptualize this, we propose a three-layer framework, as shown in Figure 1. The three layers include the input layer, the consolidation layer, and the generation layer. The input layer is the input documents of different types: official reviews and multi-turn discussions. The consolidation layer represents how meta-reviewers process the documents: they first identify and extract judgements from different documents, reorganize the judgements based on review facets, and then consolidate the opinions to form the final opinions of each review facet. In the generation layer, the meta-reviewer writes the meta-review to express the final opinions that they have developed from the previous layer.
@@ 3.2 Judgement Identification and Extraction
Judgements lay the foundation of our proposed framework and the whole peer review process. A judgement here expresses sentiment on a review facet and it contains several components: Content Expression, Sentiment Expression, Review Facet, Sentiment Level, and Convincingness Level (definitions are shown in Table 1, and an example is given in Appendix Figure 5). To automate judgement identification and extraction, we first conduct human annotation, and then leverage in-context learning of LLMs to perform more (automatic) annotation.
@@
In human annotation, there are three types of documents including meta-reviews, official reviews, and discussions (the same definition used in Li et al. (2023a)) to be annotated. We recruit two annotators [^7] The two annotators are senior PhD students who are familiar with the peer-review process. to do this annotation (annotation instructions and design are detailed in Appendix B). 30 samples (i.e., one sample = one meta-review and its corresponding reviews and discussions) are annotated [^8] Annotating one sample takes about one hour on average and it costs about 60 hours and 2,100 US dollars in total., and in total, we have 1,812 and 1,744 judgements from the two annotators. The statistics of these 30 samples are presented in Table 2. We present the agreement of the two annotators in Figure 2.999Calculation details and more results in terms of both Cohen’s $\kappa$ and Krippendorff’s $\alpha$ are in Table 11, Table 12 and Table 13 in Appendix C. Generally, we see a moderate to high agreement, suggesting that the annotation task is robust and reproducible.
@@
Figure 2: Inter-annotator agreement on meta-reviews, official reviews and discussions in terms of Krippendorff’s $\alpha$ for different judgement components including Content Expression (CE), Sentiment Expression (SE), Review Facet (RF), Sentiment Level (SL), and Convincingness Level (CL).
@@
Figure 3: The averaged GPT-4’s agreement with two human annotators on meta-reviews, official reviews and discussions in terms of Krippendorff’s $\alpha$ for different judgement components including Content Expression (CE), Sentiment Expression (SE), Review Facet (RF), Sentiment Level (SL), and Convincingness Level (CL).
To get more annotated judgements for further experiments and analysis and investigate whether LLMs can be prompted to identify and extract judgements, we split the annotation task into two sub-tasks, extracting content and sentiment expressions and predicting other components of judgements, and use GPT-4 (OpenAI, 2023) with in-context learning (see full prompts in Appendix D and E respectively for the two sub-tasks). [^10] The version of GPT-4 we use is gpt-4-0613. We present the average agreement of GPT-4 with the two human annotators in Figure 3.111111More agreement results are in Table 14, Table 15 and Table 16 in Appendix C. We can see GPT-4 has a moderate agreement with human annotators for meta-reviews and official reviews, but a low agreement for discussions. We suspect this may be because the discussions often contain rebuttals which have a different language to reviews or meta-reviews and extracting judgements from them may be more difficult. Interestingly, we also see that GPT-4 has a poor agreement in terms of convincingness (Figure 3), although the human inter-annotator agreement isn’t strong in the first place (Figure 2). These observations suggest convincingness is perhaps a subjective assessment.
@@ 3.3 Sentiment Fusion for Consolidation
Table 3: Frequency of different review facets in meta-review judgements and meta-review documents.
With all the annotated judgements extracted by humans and GPT-4, we next dive more into the process of sentiment aggregation. Among all the review facets, we find that Soundness and Advancement are the two most important review facets when the meta-reviewers write their meta-reviews, while Compliance is rarely an issue in meta-reviews (shown in Table 3). This is consistent with our understanding of the peer-reviewing process.
More importantly, we find that human meta-reviewers do not always follow the majority review sentiment. We find that in PeerSum there are 23.7% samples where the meta-reviewer’s acceptance decision is not consistent with the prediction based on majority voting by review ratings (a sample is defined as consistent when the number of reviews whose rating $\geq 5$ is larger than the number of reviews whose rating $<5$ and the final decision is Accept). We present an example in Table 4 where the meta-review does not follow the majority view on Novelty from the reviews.
@@
To understand how well the judgements from the source documents (i.e., reviews and discussions) predict the overall sentiments in the meta-reviews for each review facet, we next formulate a text classification task where the output is the sentiment level of a content expression for a review facet in the meta-review, and the input is either: (1) the annotated judgements for the facet from the source documents; or (2) the full text of the source documents. We (zero-shot) prompt GPT-4 (full prompt detailed in Appendix F) with either input to predict 100 randomly sampled human-annotated instances and present the results in Table 5. Using judgements only as input, we see that it works better in 4 out of 6 facets — this preliminary result suggests our framework of extracting these judgements as an intermediate step may help generate better meta-reviews.
Review Facets Judgements Full Texts Advancement 0.677 0.697 Soundness 0.684 0.667 Novelty 0.700 0.650 Overall 0.643 0.631 Clarity 0.712 0.645 Compliance 0.555 0.593
Table 5: Accuracy of GPT-4 in predicting the sentiment levels in meta-reviews for each facet, using either only the annotated judgements (“Judgements”) or the full text (“Full Texts”) from the source documents.